A feather falls to the earth at the same time a gust of wind pushes the feather sideways with a force of 10.2 n If the mass of the feather is 0.323 kg, what is the acceleration of the feather?


1
Expert's answer
2021-02-22T16:03:21-0500
"a=\frac{F}{m}\\a=\frac{10.2}{0.323}=31.6\frac{m}{s^2}"